solve for the variable in each equation below. show all work please. check the answer. if solution is undefined, state that in the answer. thank you for your help.
(5a-3/2)-(3a-9/8)=6

Answer by stanbon(57984)

(5a-3)/2 -(3a-9)/8 = 6
----
The least common multiple of 2 and 8 is 8.
Multiply thru by 8 to get:
4(5a-3) - (3a-9) = 8*6
---
20a-12-3a+9 = 48
---
17a - 3 = 48
---
17a = 51
---
a = 3
===========
Cheers,
Stan H.